The specific heat capacity of copper metal is 0.38 J /K.g. Calculate 

the temperature rise of a 100 g bar of copper when 250 J of heat are 

transferred to it.



1
Expert's answer
2021-02-17T10:56:19-0500
"Q=mc\Delta T,""\Delta T=\dfrac{Q}{mc}=\dfrac{250\ J}{0.1\ kg\cdot380\ \dfrac{J}{kg\cdot \!^{\circ}C }}=6.58\ ^{\circ}C."